# A Comprehensive Survey on Data Augmentation

### v3

@@ Abstract
Data augmentation is a series of techniques that generate high-quality artificial data by manipulating existing data samples. By leveraging data augmentation techniques, AI models can achieve significantly improved applicability in tasks involving scarce or imbalanced datasets, thereby substantially enhancing AI models’ generalization capabilities. Existing literature surveys only focus on a certain type of specific modality data, and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process. To bridge this gap, this survey proposes a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities by investigating how to take advantage of the intrinsic relationship between and within instances. Additionally, it categorizes data augmentation methods across five data modalities through a unified inductive approach.
Index Terms:
Data augmentation, Data-centric taxonomy, Multi-modality
@@ I Introduction
With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition [1] and text translation [2] to complicated scientific research tasks such as drug-drug interaction prediction [3] and single-cell RNA sequencing analysis [4]. Recently, powered by Latent Diffusion Models [5] and Transformer [6], Stable Diffusion [5] and ChatGPT [7] have greatly changed human beings’ working, living, and entertainment. The success of AI products is usually attributed to AI models’ in-depth understanding of the accumulated data, thereby intrinsically uncovering data patterns and learning data-task correlations. The performances of AI models will be affected by the quantity and quality of training data, e.g. models trained with limited data suffer from over-fitting with significantly degenerative performance on testing datasets, and models trained with imbalanced samples will lead to poor generalization ability. In most situations, researchers have to overcome difficulties caused by data scarcity and distribution imbalance. An intuitive solution to these problems is to acquire more data, but data are limited or difficult to collect in many cases, and labeling the data is yet another labor-intensive work.
@@
To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient [8]. The core idea of data augmentation is to artificially enlarge the training dataset by creating modified copies of existing data. It also introduces more diversity and fills the gap between training datasets and real-world applications. As technology continues to evolve, a variety of techniques are applied to augment data samples. Some simple approaches only randomly mask part of the data [9, 10, 11], while more sophisticated augmentation schemes involve generative adversarial networks [12] or reinforcement learning agents [13].
@@
An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and thus provide insight into the nature of data augmentation. Existing literature surveys summarize data augmentation methods from different perspectives [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. As summarized in Table. I, each of them concentrates on a certain data modality. Most of them categorize data augmentation methods with modality-specific (e.g. image modality and text modality) or operation-centric (e.g. feature scaling operation, data perturbation operation) taxonomies. These surveys reflect the development of data augmentation and its application in different learning scenarios. However, they fail to cover data augmentation methods across data types, uncover their common patterns despite different modalities, and thus limit readers’ understanding of the essence of data augmentation. To fill this gap, we investigate data augmentation methods comprehensively, across the five most popular data modalities. We analyze how data augmentation methods utilize different sample numbers and how they leverage different components of the data information. Then, we propose a unified taxonomy that focuses on the way data augmentation leverages information from samples. Finally, we investigate up-to-date literature and categorize them by our proposed taxonomy.
@@
  • We propose a novel modality-independent taxonomy from a data-centric perspective that accommodates data augmentation techniques for all modalities consistently and inductively;
  • To the best of our knowledge, this is the first survey that covers data augmentation techniques across five data modalities, i.e. image, text, graph, tabular, and time-series modalities;
  • We investigate how information is consistently contained in each of these data modalities and can be utilized for data augmentation;
  • We include and categorize up-to-date literature in data augmentation.
@@ II-A Early Evolution
Despite the fact that data augmentation only became popular in no more than a decade, some of its embryonic forms were proposed quite early. The concept of data augmentation was conceived at the very beginning of Deep Learning. For example, the use of random distortion can be found in LeNet [33], where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error. Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) [34, 35]. AlexNet [36] explicitly employs several data augmentation techniques to reduce overfitting. It augments the datasets by extracting patches from images and altering the color intensity. It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaption. Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods [9]. These applications of data augmentation mainly focus on increasing the size of the training dataset and introducing diversity to it. SMOTE [37] is another early and representative work on data augmentation but from a different perspective. Focusing on addressing the class imbalance problem, it suggests that oversampling the minority class can achieve better classification performance when categories are not equally represented in the dataset.
@@ II-B Existing Surveys
Data augmentation on text data has not been thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components. Still, several augmentation methods are proposed for or transplanted to text data. [18] is among the first surveys on text data augmentation. It adopts a simple taxonomy but gives more emphasis on the applications. Built on previous work on image data augmentation, [19] continues to investigate how data augmentation is used to enrich text datasets, and methods are classified as symbolic augmentation and neural augmentation. [20] categorizes text data augmentation methods by identifying whether they are used on the data space or feature space. It also presents the performance improvements of different methods. [31] focuses on the application of data augmentation for training Large Language Models. [32] discusses how Multimodal Large Language Models can be used for augmenting text, image, and speech data.
@@
Till now, data augmentation techniques for different data types have been exhaustively discussed; however, there has not been a comprehensive survey that summarizes data augmentation for all types of data. Such a survey can be profitable because data augmentation techniques for different modalities share some common methodologies when leveraging information from existing data, such as altering some features of a sample or mixing values between multiple samples. Summarizing these similarities can help reveal the common pattern of data augmentation. A comprehensive survey does not mean simply combining existing surveys because the proposed taxonomies in these surveys view data augmentation from different angles and are incompatible with each other. Some taxonomies contrast basic operations to advanced approaches; some categorize methods with modality-specific terms. Besides, these taxonomies provide limited insight into what happens to the information. Indeed, huge gaps lie between operations for data modalities due to their distinctive nature; these operation-based or modality-specific taxonomies reflect the development of data augmentation techniques. However, from such taxonomies, we cannot grasp the essence of data augmentation and lack perception of the big picture of data augmentation. Hence there is a need for a comprehensive survey that covers data augmentation techniques for all data modalities with an all-embracing and data-centric taxonomy.
@@ III Taxonomy
$f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ (1)
@@
In this survey, we try to propose a two-tier taxonomy from a data-centric perspective that can be applied to all data modalities. The main consideration of our taxonomy is whence the information in the augmented data comes. It differentiates data augmentation techniques by asking two research questions: RQ1: How many samples are used to generate each new sample? RQ2: Which part of the information is used to generate new data? The answers to these questions constitute the taxonomy hierarchy which we present below.
@@
$\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}$ (2)
@@
$\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}$ (3)
@@
Dataset Level Augmentation. This type of data augmentation method does not explicitly make use of one or multiple data but rather uses the entire dataset. It comprehensively learns the features and abstractions of the dataset, and samples completely new data from a learned distribution of the dataset. The generated data should also fall into this distribution. We formalize dataset level augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
$\vspace{-1.5mm}\mathbf{\tilde{x}}\sim P(\mathbf{X}),\qquad\tilde{y}\sim P(\mathbf{y}|\mathbf{\tilde{x}})$ (4)
@@
We have presented a general taxonomy for data augmentation at a high level with abstract and conceptual formalization. We organize this survey in accordance with this taxonomy, from Section IV to Section VI. We elaborate on how the proposed taxonomy applies to each data modality and review some typical data augmentation methods.
@@ IV Single-instance Level Augmentation
For image data, the elemental components are pixels. A pixel carries unit information that uses a numerical value to represent the color. Each pixel also has spatial relationships with other pixels, so, images are well-structured. Single-instance level augmentation for image data revolves around the perturbation of the image pixels’ colors and spatial relationships. For text data, the elemental components of text data are words (or sometimes, tokens). A word often expresses minimal meanings. Words are organized into sentences following syntactic rules. The sentence is a sequential representation since one word comes after another, but there is also a tree-structured grammatical dependency behind the sentence. Single-instance level augmentation for text data often changes the wording or reshapes the syntax of a sentence. A graph is represented by $\mathbf{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ stands for the vertex set, and $\mathcal{E}$ stands for the edge set. The topology structure can be alternatively represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}$ is $1$ if there is an edge connecting vertex $i$ and vertex $j$, and is 00 otherwise. In a deep graph learning scenario, the graph also contains a feature matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ that represents the attributes of the nodes. Hence another representation for a graph is $\mathbf{G}=\{\mathbf{A},\mathbf{X}\}$. Single-instance level augmentation for graph data is to perturb the values in the feature matrix $\mathbf{X}$ of nodes or the topology structure represented by the adjacency matrix $\mathbf{A}$. A table is made up of cells, which store numerical or categorical values that represent attributes of the described entities. Cells are organized into columns and rows, but this structural information is trivial, and switching the sequence of the columns and rows does not affect the information conveyed by the table. So, most data augmentation methods leverage the value information. Still, it is also possible to perturb the structure information of a table. Time-series data represent sequential values within a period, and each timestamp is considered an elemental component. Data augmentation methods for time-series data either perturb the value at each timestamp or perturb the sequence of timestamps.

### v4

@@ Abstract
Data augmentation is a series of techniques that generate high-quality artificial data by manipulating existing data samples. By leveraging data augmentation techniques, AI models can achieve significantly improved applicability in tasks involving scarce or imbalanced datasets, thereby substantially enhancing AI models’ generalization capabilities. Existing literature surveys only focus on a certain type of specific modality data and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process. To bridge this gap, this survey proposes a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities by investigating how to take advantage of the intrinsic relationship between and within instances. Additionally, it categorizes data augmentation methods across five data modalities through a unified inductive approach.
@@ I Introduction
With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition [1] and text translation [2] to complicated scientific research tasks such as drug-drug interaction prediction [3] and single-cell RNA sequencing analysis [4]. Recently, Stable Diffusion [5] and ChatGPT [6] have greatly changed human beings’ working, living, and entertainment. The success of AI products is usually attributed to AI models’ in-depth understanding of the accumulated data, thereby intrinsically uncovering data patterns and learning data-task correlations. The performances of AI models will be affected by the quantity and quality of training data, e.g., models trained with limited data suffer from overfitting with significantly degenerative performance on testing datasets, and models trained with imbalanced samples will lead to poor generalization ability. In most situations, researchers have to overcome difficulties caused by data scarcity and distribution imbalance. An intuitive solution to these problems is to acquire more data, but data are limited or difficult to collect in many cases, and labeling the data is yet another labor-intensive task.
@@
To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient [7]. The core idea of data augmentation is to artificially enlarge the training dataset by creating modified copies of existing data. It also introduces more diversity and fills the gap between training datasets and real-world applications. Data augmentation is widely adopted in the real world and exhibits success in both the dataset preparation stage, such as the ImageNet-C and ImageNet-P datasets [8], which use data augmentation to simulate common image corruptions and perturbations, as well as the model training stage, such as scSiameseClu [9], which uses data augmentation to generate dual views for graph representation learning. As technology continues to evolve, a variety of techniques are applied to augment data samples. Some simple approaches only randomly mask part of the data [10, 11, 12], while more sophisticated augmentation schemes involve generative adversarial networks [13] or reinforcement learning agents [14].
@@
An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and provide insight into the nature of data augmentation. Existing literature surveys summarize data augmentation methods from different perspectives [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. As summarized in Table I, each survey concentrates on a certain data modality. Most of them categorize data augmentation methods with modality-specific (e.g., image modality and text modality) or operation-centric (e.g., feature scaling operation and data perturbation operation) taxonomies. These surveys reflect the development of data augmentation and its application in different learning scenarios. However, they fail to cover data augmentation methods across data types, uncover their common patterns despite different modalities, and thus limit readers’ understanding of the essence of data augmentation. For example, originating from the idea of interpolation, mixup-based augmentation has been applied across different modalities, whereas such consistency is ignored in modality-specific discussions. To fill this gap, we investigate data augmentation methods comprehensively across the five most popular data modalities. We analyze how data augmentation methods utilize different sample numbers and how they leverage different components of the data information. Then, we propose a unified taxonomy that focuses on the way data augmentation leverages information from samples. Finally, we investigate up-to-date literature and categorize it by our proposed taxonomy.
@@
$\bullet$ We propose a novel modality-independent taxonomy from a data-centric perspective that accommodates data augmentation techniques for all modalities consistently and inductively.
$\bullet$ To the best of our knowledge, this is the first survey that covers data augmentation techniques across five modalities: image, text, graph, tabular, and time-series data.
$\bullet$ We investigate how information is consistently contained in each of these modalities and can be utilized for augmentation.
$\bullet$ We include and categorize up-to-date literature in data augmentation and discuss future directions.
@@ II-A Early Evolution
Despite the fact that data augmentation only became popular in no more than a decade, some of its embryonic forms were proposed quite early. The concept of data augmentation was conceived at the very beginning of deep learning. For example, the use of random distortion can be found in LeNet [34], where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error. Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) [35]. AlexNet [36] explicitly employs data augmentation to reduce overfitting. It augments the datasets by extracting patches from images and altering the color intensity. It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaptation. Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods [10]. These applications of data augmentation mainly focus on increasing the size of the training dataset and introducing diversity to it. SMOTE [37] is designed from a different perspective. Focusing on addressing the class imbalance problem, it suggests that oversampling the minority class can achieve better classification performance when categories are not equally represented in the dataset.
@@ II-B Existing Surveys
Data augmentation on text data has not been thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components. Still, several augmentation methods are proposed for text data. [19] is among the first surveys on text data augmentation. It adopts a simple taxonomy but gives more emphasis on the applications. Built on previous work on image data augmentation, [20] continues to investigate how data augmentation enriches text datasets, and methods are classified as symbolic augmentation and neural augmentation. [21] categorizes text data augmentation methods by identifying whether they are used on the data space or feature space. It also presents the performance improvements of different methods. [32] focuses on the application of data augmentation for training large language models. [33] discusses how multimodal large language models can be used for augmenting text, image, and speech data.
@@
Till now, data augmentation techniques for different data types have been exhaustively discussed; however, there has not been a comprehensive survey that summarizes data augmentation for all types of data. Such a survey is profitable because data augmentation techniques for different modalities share some common methodologies when leveraging information from existing data, such as altering some features of a sample or mixing values between multiple samples. Summarizing these similarities can help reveal the common pattern of data augmentation. A comprehensive survey does not mean simply combining existing surveys because the proposed taxonomies in these surveys view data augmentation from different angles and are incompatible with each other. Some taxonomies contrast basic operations to advanced approaches; some categorize methods with modality-specific terms. Besides, these taxonomies provide limited insight into what happens to the information. Indeed, huge gaps lie between operations for data modalities due to their distinctive nature; these operation-based or modality-specific taxonomies reflect the development of data augmentation techniques. However, from such taxonomies, we cannot grasp the essence of data augmentation and lack perception of the big picture of data augmentation. Hence there is a need for a comprehensive survey that covers data augmentation techniques for all data modalities with an all-embracing and data-centric taxonomy.
@@ III Taxonomy
$f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}\vskip-5.69054pt$ (1)
@@
In this survey, we try to propose a two-tier taxonomy from a data-centric perspective that can be applied to all data modalities. The main consideration of our taxonomy is whence the information in the augmented data comes. It differentiates data augmentation techniques by asking two research questions: RQ1: How many samples are used to generate each new sample? RQ2: Which part of the information is used to generate new data? The answers to these questions constitute the taxonomy hierarchy, which we present below.
@@
$\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}\vskip-5.69054pt$ (2)
@@
$\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}\vskip-5.69054pt$ (3)
@@
Dataset Level Augmentation. This type of data augmentation method does not explicitly make use of one or multiple data points but rather uses the entire dataset. It comprehensively learns the features and abstractions of the dataset and samples completely new data from a learned distribution of the dataset. The generated data should also fall into this distribution. We formalize dataset level augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
[…]
@@
We have presented a general taxonomy for data augmentation at a high level with abstract and conceptual formalization. We organize this survey in accordance with this taxonomy, from Section IV to Section VI. We elaborate on how the proposed taxonomy applies to each data modality and review some typical data augmentation methods.  [^1] A collection of discussed papers is available at https://github.com/ZaitianWang/Awesome-Data-Augmentation.
TABLE II: Comparative summary of all methods, including detailed categories, operations, and corresponding analyses.
# of samples Information type Modality Category Operation Analysis Computation Info loss
Single-instance Value Image Pixel erasing Applying a rectangular- or grid-shaped mask over the image Moderate Negligible High
Photometric transformation Brightness adjustment, color inversion, color casting, greyscaling, histogram equalization, kernel filters Negligible Moderate Negligible
Text Token replacement Synonym, embedding similarity, knowledge graph Negligible Moderate Negligible
Token addition Synonym, random punctuation Negligible Negligible Negligible
Token deletion Randomly dropping tokens Negligible Negligible High
Graph Node attribute masking Applying a mask over the node attribute matrix to hide certain nodes or features Negligible Negligible High
Tabular Table masking Applying a mask over the table to hide some cells Negligible Negligible High
Feature engineering Arithmetic operations High High Negligible
Time-series Amplitude perturbation Scaling, magnitude warping, rotation, and jittering Negligible Negligible Negligible
Sequence decomposition Extract components from a sequence High Moderate High
Structure Image Image cropping Selecting an area and removing other parts Moderate Negligible High
Geometric transformation Flipping, shearing, rotation, translation, stretching, distortion Negligible Moderate Negligible
Text Sentence cropping Dividing sentence to pieces Moderate Negligible High
Sentence morphing Inversion, passivization Moderate Negligible Negligible
Graph Topology perturbation Node addition, edge addition, node dropping, edge dropping Negligible Negligible Moderate
Subgraphing Cropping or substituting subgraphs Moderate Negligible High
Graph diffusion Transition matrices sampled from random walk High Moderate Moderate
Graph rewiring Swapping edges High Moderate Moderate
Tabular Table subsetting Splitting table by columns Negligible Negligible High
Time-series Window Slicing Selecting a span from a sequence Negligible Negligible High
Window Morphing Downsampling, upsampling, shuffling Negligible Negligible Negligible
Value-structure Image Policy-based DA Selecting an group of augmentation operations with an RL agent High High Moderate
Prompt-based image editing Prompting visual generative models to edit image Negligible High Negligible
Text Hierarchical data augmentation Augmenting at both word and sentence level High Negligible Moderate
Back-translation Translating to an intermediate language and translating back to the original language Negligible High Negligible
Paraphrase Rewriting sentence with alternative expressions while conveying the same meanings Negligible High Negligible
Multi-instance Value Image Image mixup Mixing color from multiple images Negligible Negligible Moderate
Text Text mixup Mixing word or sentence embeddings Negligible Moderate Moderate
Graph Graph propagation Spreading node features or labels through edges to adjcacent nodes Moderate Negligible Moderate
Tabular Tabular mixup Mixing cell values across rows Negligible Moderate Moderate
Time-series Sequence averaging Calculating averaged sequence values Negligible Moderate Moderate
Structure Image Image Patching Putting image segments together Negligible Negligible High
Text Text fragment merging Swapping substructures between sentences Moderate Negligible High
Tabular Table Concatenation Unifying table columns to concatenate tables by overlapping columns Moderate Negligible Negligible
Value-structure Image Neural blending Mixing images through a neural network Negligible High Negligible
Prompt-based mixing Prompting Diffusion Models to combine images Negligible High Negligible
Text Generative mixing Prompting LLMs to mix sentences Negligible High Negligible
Graph Graph mixup Aligning the topology and mixing node features Moderate Moderate Negligible
Dataset Value-structure Cross-modality AE Using reconstructed sample as augmentation Negligible High Negligible
GAN Using the GAN’s generator for data augmentation Negligible High Negligible
LLM Prompting the LLM to generate large amount of textual samples Negligible High Negligible
Diffusion Using discrete denoising diffusion model to generate new graphs Negligible High Negligible
Exogenous Image, text Neural Style Transfer Introducing styles from other datasets High Moderate Negligible
Tabular Table Retrieval Retrieve external database to fill and expend existing tables High Negligible Negligible
Image Computer Graphic Modeling Designing scenes and objects with computer softwares High High Negligible
Tabular Relational Structure Construction Building external graph structures over the table High Moderate Negligible
Time-series Statistical Generation Using expert knowledge to expand sequences High Moderate Negligible
@@ IV Single-instance Level Augmentation
For image data, the elemental components are pixels. A pixel carries unit information that uses a numerical value to represent the color. It also has spatial relationships with other pixels. Single-instance level image augmentation perturbs the image pixels’ colors or spatial relationships. For text data, the elemental components are words or tokens. A word often expresses minimal meanings. Words are organized into sentences following syntactic rules. Single-instance level text augmentation often changes the wording or reshapes the syntax of a sentence. A graph is represented by $\mathbf{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ and $\mathcal{E}$ stand for the vertex set and the edge set. The topology structure can be alternatively represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$. The feature matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ represents the attributes of the nodes. Hence another representation for a graph is $\mathbf{G}=\{\mathbf{A},\mathbf{X}\}$. Single-instance level graph augmentation perturbs the values in the feature matrix $\mathbf{X}$ or the topology structure represented by the adjacency matrix $\mathbf{A}$. A table is made up of cells with numerical or categorical values that represent attributes of the described entities. Cells are organized into columns and rows, but this structural information is trivial, and switching the sequence of the columns and rows does not affect the information conveyed by the table. So, most data augmentation methods leverage the value information with only a few exceptions. Time-series data represent sequential values within a time span, and each timestamp is considered an elemental component. Data augmentation methods for time-series data either perturb the value at each timestamp or perturb the sequence of timestamps.